OpenStudy (anonymous):

Without substitution, solve the formula P=2L+2W for the unknown variable W in terms of the known variables P,L. Enter the rewritten equation.

OpenStudy (campbell_st):

W = (P -2L)/2

OpenStudy (anonymous):

p = 2l+2w implies p = 2(l+w) implies l+w = p/2 implies w = p/2 - l